# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

### v1

@@ Acknowledgements
Appendix A Appendix
Contents
@@ A.1 Algorithm of Autotune Lasso with active set selection
Input: Predictors xx(normalized), response $y$, Sequential F-test significance level $\alpha$
@@
21exInitialize $\widehat{\beta}\leftarrow 0$, $r\leftarrow y$, $\widehat{\sigma}^{2}\leftarrow\mathrm{Var}(y)$, $\lambda_{0}\leftarrow\frac{1}{\mathrm{Var}(y)}\|\frac{x^{\top}y}{2n}\|_{\infty}$, $\mathsf{Predictor.Ranking}\leftarrow 1:p$, $\mathsf{Support.Set}\leftarrow\phi$, sigma.update.flag $\leftarrow\mathsf{TRUE}$
@@
9    [2pt] $\widehat{\beta}_{j}\leftarrow\mathsf{Soft.Threshold}_{\lambda_{0}\widehat{\sigma}^{2}}\mathopen{}\mathclose{{\left(\frac{1}{n}x_{\cdot,j}^{\top}r+\widehat{\beta}_{j}}}\right)$
10    [2pt] $r\leftarrow r+x_{\cdot,j}\mathopen{}\mathclose{{\left(\widehat{\beta}^{\mathsf{(old)}}_{j}-\widehat{\beta}_{j}}}\right)$
@@
12 Calculate partial residuals $r^{(i)}\leftarrow r+x_{\cdot,i}\widehat{\beta}_{i}$
@@
28 $\mathrm{error}\leftarrow\frac{\|\widehat{\beta}-\widehat{\beta}^{\mathsf{(old)}}\|_{1}}{\max(\|\widehat{\beta}^{\mathsf{(old)}}\|_{1},10^{-8})}$
@@
Output: regression coefficients’ estimate $\widehat{\beta}$
@@
5 [2pt] if $\mathopen{}\mathclose{{\left|\frac{1}{n}\langle{x_{\cdot,j}},{r}\rangle}}\right|\geq\lambda$ then
6     Add $x_{j}$ to the back of $\mathsf{Active.Set}$
@@
9     $r\leftarrow r+x_{j}\widehat{\beta}_{j}$
@@ A.2.4 Accuracy when only approximate sparsity holds
Table 5: Table of mean runtimes (with SD in the brackets) for Fig.˜7(b). $n=200,s=15,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 100 to 1000.
@@ A.6 Supplementary plots for Data Analysis Section
(b) train set size $n=170$, test set size 252n=82252-n=82.
(c) train set size $n=180$, test set size 252n=72252-n=72.
(d) train set size $n=200$, test set size 252n=52252-n=52.
(e) train set size $n=210$, test set size 252n=42252-n=42
@@
(b) train set size $n=170$, test set size 252n=82252-n=82.
(c) train set size $n=180$, test set size 252n=72252-n=72.
(d) train set size $n=200$, test set size 252n=52252-n=52.
(e) train set size $n=210$, test set size 252n=42252-n=42

### v2

@@ Acknowledgements
Table of Contents for the Appendix
Appendix A Supplementary algorithms and simulation results
We report additional algorithms and simulation results in this section.
@@ A.1 Algorithm of Autotune Lasso with active set selection
Input: Predictors X (normalized), response Y, sequential F-test significance level $\alpha$
@@
21exInitialize $\widehat{\beta}\leftarrow 0$, $r\leftarrowY$, $\widehat{\sigma}^{2}\leftarrow\mathrm{Var}$(Y), $\lambda_{0}\leftarrow\frac{1}{\mathrm{Var}(\mathrm{Y})}\|\frac{\mathrm{X}^{\top}\mathrm{Y}}{2n}\|_{\infty}$, $\mathsf{Predictor.Ranking}\leftarrow 1:p$, $\mathsf{Support.Set}\leftarrow\phi$, sigma.update.flag $\leftarrow\mathsf{TRUE}$
@@
9    [2pt] $\widehat{\beta}_{j}\leftarrow\mathsf{Soft.Threshold}_{\lambda_{0}\widehat{\sigma}^{2}}\mathopen{}\mathclose{{\left(\frac{1}{n}\mathrm{X}_{\cdot,j}^{\top}r+\widehat{\beta}_{j}}}\right)$
10    [2pt] $r\leftarrow r+\mathrm{X}_{\cdot,j}\mathopen{}\mathclose{{\left(\widehat{\beta}^{\mathsf{(old)}}_{j}-\widehat{\beta}_{j}}}\right)$
@@
12 Calculate partial residuals $r^{(i)}\leftarrow r+\mathrm{X}_{\cdot,i}\widehat{\beta}_{i}$
@@
28 $\mathrm{error}\leftarrow\|\widehat{\beta}-\widehat{\beta}^{\mathsf{(old)}}\|_{1}/\|\widehat{\beta}^{\mathsf{(old)}}\|_{1}$
@@
Output: $\widehat{\beta}$
@@
5 [2pt] if $\mathopen{}\mathclose{{\left|\frac{1}{n}\langle{\mathrm{X}_{\cdot,j}},{r}\rangle}}\right|\geq\lambda$ then
6     Add $\mathrm{X}_{\cdot,j}$ to the back of $\mathsf{Active.Set}$
@@
9     $r\leftarrow r+\mathrm{X}_{\cdot,j}\widehat{\beta}_{j}$
@@ A.2.1 High Dimensional Setting: n = 80, p = 750, s = 5
Figure 14: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV Lasso, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
@@ A.2.4 Accuracy when only approximate sparsity holds
Table 5: Table of mean runtimes (with SD in the brackets) for Fig.˜6(b). $n=200,s=15,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 100 to 1000.
@@ A.6 Supplementary plots for Data Analysis Section
(b) train set size $n=160$, test set size 252n=92252-n=92.
(c) train set size $n=170$, test set size 252n=82252-n=82.
(d) train set size $n=180$, test set size 252n=72252-n=72.
(e) train set size $n=200$, test set size 252n=52252-n=52.
(f) train set size $n=210$, test set size 252n=42252-n=42
@@
(b) train set size $n=160$, test set size 252n=92252-n=92.
(c) train set size $n=170$, test set size 252n=82252-n=82.
(d) train set size $n=180$, test set size 252n=72252-n=72.
(e) train set size $n=200$, test set size 252n=52252-n=52.
(f) train set size $n=210$, test set size 252n=42252-n=42